OpenStudy (anonymous):

Sara recorded the number of points she scored in each of the 30 basketball games in one season. Which statement is best supported by this histogram? A. Sara usually scored between 20 and 29 points in a basketball game. B. Sara scored a total of 30 points this season. C. Sara scored more than 9 points in 20 games. D. Sarah didn't score in 3 games.

OpenStudy (anonymous):

0-9 = 10 10-19 = 15 20-29 = 5

OpenStudy (shrutipande9):

can u show d histogram?

OpenStudy (anonymous):

its 0 to 9 = 10 10 to 19 = 15 20 to 29 == 5

OpenStudy (anonymous):

i was thinking it may be c

OpenStudy (anonymous):

and i was right

OpenStudy (shrutipande9):

oh great..and srry i cud nt be of any help..:(

OpenStudy (anonymous):

its ok